"An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 655 hours and a standard deviation of 21 hours. Random samples of 43 bulbs are drawn from this population, and the mean of each sample is determined. What is the standard error of the mean of the sampling distribution?"


1
Expert's answer
2022-01-31T16:13:15-0500

"\mu=655\\\sigma=21\\n=43"

The standard error of the mean of the sampling distribution is given as,

"SE(\bar x)={\sigma\over\sqrt{n}}={21\over\sqrt{43}}={21\over6.55743852}=3.2025(4dp)"

Therefore, the standard error of the mean of the sampling distribution is 3.2025